What is a number whose product with 6 is the same as its sum with 60? We don't want to just know the answer, we want to know how you did it too.
Answer by rapaljer(4671)

Let x= the number

6x = x+ 60

Subtract x from each side:
6x -x = x-x+60
5x=60
x=60/5=12

Check: Take the product of 12 with 6, which is 72. This is indeed the same as its sum with 60, which is also 72.

R^2

Dr. Robert J. Rapalje